### v2

# methodsReferences (Methods): Selection bias obfuscates the discovery of fast radio burst sources

Abstract
Fast Radio Bursts (FRBs) are a newly discovered class of extragalactic radio transients characterised by their high energy and short duration (sms)[1]. Their elusive physical origin remains a subject of ongoing research, with magnetars emerging as leading candidates[2, 3]. Previous studies have employed various methodologies to address the FRB origin problem, including demographic analyses of FRB host galaxies and their local environments[4, 5, 6], assessments of FRB rate evolution with redshift[7, 8, 9], and searches for proposed multi-messenger FRB counterparts[10]. However, these studies are susceptible to significant biases stemming from unaccounted radio and optical selection effects. Here we present empirical evidence for a substantial selection bias against detecting FRBs in galaxies with large inclination angles (edge-on) using a sample of hosts identified for FRBs discovered by untargeted surveys. This inclination-related bias likely leads to a significant underestimation (by about a factor of two) of the FRB rates reported in the literature and disfavours globular clusters as the dominant origin of FRB sources, as previously speculated[6]. These conclusions have important implications for FRB progenitor models and targeted FRB follow-up strategies. We further investigate the impact of this bias on the relative rate of FRBs in different host environments. Our analysis suggests that scattering in FRB hosts is likely responsible for the observed bias[11, 12]. However, a larger sample of localised FRBs is required to robustly quantify its contribution in the inclination-related selection bias.
McWilliams Center for Cosmology, Department of Physics, Carnegie Mellon University, Pittsburgh, PA 15213, USA
†These authors contributed equally to this work
In this study, we investigate potential biases introduced by the inclination angle () of FRB hosts on the detectability of FRBs in various radio surveys. We select published FRBs localised with arcsecond precision and associated with their hosts with a PATH (Probabilistic Association of Transients to their Hosts)[13] posterior probability 95%. To minimise biases from targeted surveys, we focus on hosts identified from two untargeted surveys: the Commensal Real-time ASKAP Fast Transients Survey (CRAFT)[14] and Deep Synoptic Array (DSA)-110[15].
Our study is limited to non-repeating FRBs, as most repeating FRBs have been identified by the Canadian Hydrogen Intensity Mapping Experiment (CHIME)[16], and their precise localization ( few arcseconds) through targeted follow-up campaigns excludes them from our analysis. We also restrict our sample to hosts with an apparent r-band magnitude (mr) 21 AB to ensure robust inclination angle measurements using archival imaging survey data.
The imaging surveys used in this analysis are the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) survey[17] and the DESI (Dark Energy Spectroscopic Instrument) survey[18], which are sensitive to galaxies with mr 22 AB mag and 23.5 AB mag, respectively. We prioritise DESI survey data due to its greater depth, using Pan-STARRS data only when the FRB host is outside the DESI footprint. This approach yields a sample of 23 FRB hosts, listed in Table Selection bias obfuscates the discovery of fast radio burst sources.
To estimate the inclination angles () of the 23 FRB hosts, we employ two methods: parametric profile fitting using the software AutoProf[19] and a non-parametric elliptical isophotal fitting routine from the software Photutils[20]. The estimates from both methods are reported in Table Selection bias obfuscates the discovery of fast radio burst sources, and their consistency supports the robustness of our calculations (Methods). Notably, the best-fitted Sérsic index from the first method is 1 for all 23 FRB host galaxies, indicating that these hosts are likely disk-dominated late-type galaxies[21], as also found in a recent study focusing on local Universe FRB hosts ()[6]. However, it is important to note that at least two FRBs, both repeating sources, have been localised to irregular star-forming dwarf galaxies[22, 23]. These FRB sources are situated within the interstellar medium (ISM) of their respective hosts, coinciding with regions of active star formation.
We next compare the observed distribution of for FRB hosts with that expected from a sample of randomly selected field galaxies. Our null hypothesis asserts that the inclination distribution of FRB hosts mirrors that of field galaxies. Given that galaxies are viewed at random angles and that there is no preferred orientation in the Universe[24], we expect a uniform distribution of the cosine of inclination angles () among field galaxies[25]. However, deviations from this expected uniformity are observed, which can be attributed to the limitations inherent in the methods used to estimate (Methods).
To address these limitations and ensure a fair comparison with our FRB host sample, which are disk-dominated, we derived the expected distribution of 23 randomly sampled disk-dominated late-type galaxies from the Sloan Digital Sky Survey (SDSS) DR16 catalogue[26]. To exclude early-type galaxies, we apply the following ur colour criterion: [27]. Additionally, we use the same constraint on apparent r-band magnitude (mr 21 AB mag) as employed in selecting FRB hosts. We note that these selection criteria do not affect the conclusions of our study (Methods). For the selected galaxies, we utilise morphological parameters provided by the SDSS DR16 catalogue, including the axial-ratio (), calculated by fitting an exponential disk profile to the galaxy light, to determine their .
The cumulative distribution functions (CDFs) of the values for both the randomly sampled SDSS galaxies and FRB hosts are displayed in Figure 1. To ensure a fair comparison, we employ estimates derived from the parametric profile fitting method, which aligns with the technique used in the SDSS pipeline. Our findings remain consistent even if estimates from the non-parametric elliptical isophotal fitting method are used. Additionally, to mitigate potential sampling effects, we generate 10,000 realisations of a sample comprising 23 randomly selected SDSS disk-dominated galaxies with m 21 AB mag. From these realisations, we estimate a 68% credible bound on the SDSS galaxy CDF, also illustrated in Figure 1.
To compare the distributions of the FRB host sample with randomly sampled SDSS galaxies, we use two non-parametric tests: the two-sample Anderson-Darling (AD) test[28] and the Mann-Whitney (MW) U test[29]. The p-values obtained, from the AD test and from the MW test, signify a significant difference in the distributions between the localised FRBs and the SDSS galaxy sample. Hence, we conclude that the distribution of our FRB host sample distinctly differs from that of field galaxies in the Universe.
An alternative approach to assess the difference in CDFs involves dividing the observed mean CDF of SDSS galaxies into bins. We opt for four bins, which ensures at least one FRB host galaxy per bin. Note that increasing the number of bins would result in empty bins, which would only strengthen our conclusion. These bins correspond to the quartiles of the histogram for the SDSS galaxies, ensuring a uniform distribution of among these randomly sampled galaxies.
The range for each bin is listed in Table 1. In Figure 2a, we illustrate the distribution of FRB hosts across the four bins, which does not follow a uniform distribution. The majority of FRBs (48% of the FRB hosts) have , and there is an evident scarcity of edge-on galaxies (only one with ). This inclination-related bias is purely observational, as there is no inherent reason for FRB sources to be preferentially located in galaxies appearing nearly face-on to us, suggesting a significant underestimation of FRB all-sky rates. Based on the number of FRB hosts in Bin 1 (11) and disregarding differences in exposure and sensitivity between DSA-110 and CRAFT surveys, this underestimation (44 expected instead of 23 observed) can be a factor of two assuming Poisson statistics. A more robust estimate would require a larger sample of FRB hosts and must consider the various systematic factors present in each survey.
This has two major implications: First, it reinforces previous assertions made using the reported all-sky rate in the literature, suggesting that the majority of FRBs are likely repeating sources[30, 31, 32]. Second, any targeted follow-up campaigns aiming to determine the relative rates of various FRB formation channels, such as observations of nearby star-forming galaxies[33] or galaxy clusters[34], would need to account for the noted inclination-related bias.
Furthermore, the observed dependence of the FRB rate on of the hosts suggests that the majority of FRB sources do not constitute a spherical or halo population, such as globular clusters in the halos of their hosts. If this were the case, one would expect to observe either more FRBs in inclined host bins (Bin 4 in Figure 2a) than in face-on ones (Bin 1) as a significant fraction of globular clusters are present in the inner halo close to their host disk[35], or no significant difference in FRB distribution if FRB sources are uniformly distributed in their host halo. Therefore, this suggests a disk origin for the majority of FRB sources. Consequently, FRB 20200120E, the closest extragalactic FRB localised to a globular cluster in the Messier 81 galaxy[36, 37], may be considered as an outlier in the known FRB sample.
While the current sample size of FRB hosts may not definitively explain the observed inclination-related bias, one plausible explanation could involve propagation-induced transmission effects, such as dispersion and scattering. For instance, the contribution of the host galaxy’s ISM to the FRB DM likely varies significantly with the inclination angle of the host disk[38]. In Figure 2b, we show a box-plot illustrating the range of DM values contributed by FRB hosts (DMhost) in each of the four bins. The DMhost for our sample FRB hosts are listed in Table Selection bias obfuscates the discovery of fast radio burst sources. In general, higher DMhost values are expected in bins corresponding to higher (Bins 3 and 4). However, the sample utilised in this study does not exhibit this trend, indicating that we may be overlooking FRBs with significant DMhost in Bins 3 and 4. To robustly quantify the impact of DMhost on FRB hosts in Bins 3 and 4, we require a larger number of FRB hosts across all four bins.
The potential scarcity of FRBs with large DMhost in Bins 3 and 4 can be linked to pronounced scattering in the host ISM (Methods), akin to observations seen for pulsars within the Milky Way[39]. Scattering has also been previously recognised as a major factor influencing the observed properties and detectability of FRBs[11, 12]. Furthermore, both DMhost and scattering are expected to be more prominent in star-forming galaxies, where the ISM likely contains a substantial amount of dense, ionised, and turbulent gas[40]. This can lead to a higher expected scattering (based on the observed star-formation rates (SFR)gas velocity dispersion relation[41]) for the same DM compared to a galaxy like our own Milky Way[42]. Additionally, the probability of the FRB sight-line embedded in the Galactic disk intersecting ionised clouds, such as star-forming regions and HII regions, would also be higher in star-forming galaxies. These effects are particularly severe for FRB sources located well within the disk of edge-on galaxies, resulting in a larger inclination-based bias, akin to what is proposed in the case of core-collapse supernovae[43].
To examine this hypothesis, we analyse the distribution of SFR among FRB host galaxies across the four bins. Figure 2c illustrates this distribution for 23 FRB host galaxies, while Figure 2d focuses on star-forming host galaxies (defined as those with specific SFR (sSFR) yr-1[44]). Two observations can be made from these plots: Firstly, FRB hosts with higher exhibit lower median SFRs, indicating a potential inclination-related bias. Secondly, the proportion of star-forming galaxies increases in Bins 3 and 4. This observation is noteworthy as it hints at the possibility of FRB sources having a preference for star-forming galaxies, despite the anticipated propagation effects being more pronounced in such environments. Such findings lend support to the hypothesis of young FRB progenitor channels[45]. However, the limited sample size prohibits definitive conclusions from this analysis. Nevertheless, if this trend persists in future studies, it could elucidate the significant portion of highly scattered FRBs within the CHIME/FRB sample, potentially associated more often with star-forming galaxies[46]. Unfortunately, we lack scattering parameters (e.g., scattering timescale, decorrelation bandwidth) for most FRBs in our sample to test this prediction.
We also examine the distribution of redshifts among the FRB hosts in our sample across the four bins, shown in Figure 2e. We find that all but Bin 4 is consistent with a uniform distribution of FRB redshifts among the four bins, albeit with a small sample size. Given the potential correlation between FRB host properties, such as SFR, and , any bias in the FRB redshift distribution due to host inclinations could potentially bias studies aiming to constrain the origins of FRBs, such as investigations into the evolution of FRB properties relative to SFR or stellar mass. Therefore, if the noted trend is found to be true in case of larger sample of FRB hosts, it would suggest the necessity of correcting for the inclination-related selection effect in the aforementioned analyses. Note that can also influence host demographics studies involving projected FRB offset measurements[4], which will be explored elsewhere.
In conclusion, our study highlights the importance of investigating inclination-related biases using a larger sample of both repeating and non-repeating FRBs from upcoming wide-sky sensitive surveys. To mitigate these biases, we recommend focusing on FRB hosts with in future studies.
Sample of 23 FRB host galaxies employed in this work Numbera FRB Inclination angleb Bin # SFR0-100Myrc log(sSFR) Mass z DM Host mrd Reference Degrees M⊙ yr-1 log(yr-1) M pc cm-3 AB mag 1 20200906A 63, 59 Bin3 4.93 9.43 1.33 0.3688 309 19.95 [5] 2 20210410D 41, 41 Bin1 0.03 10.99 0.30 0.1415 468 20.65 [5] 3 20211127I 21, 26 Bin1 35.83 7.93 0.30 0.0469 171 14.97 [5] 4 20211203C 7, 11 Bin1 15.91 8.56 0.58 0.3437 395 19.64 [5] 5 20211212A 33, 35 Bin1 0.73 10.42 1.91 0.0707 127 16.44 [5] 6 20220310F 39, 41 Bin1 4.25 9.34 0.93 0.4780 22 20.72 [15] 7 20220914A 45, 46 Bin2 0.72 9.62 0.30 0.1139 538 19.77 [15] 8 20220920A 27, 29 Bin1 4.1 9.2 0.65 0.1582 170 18.16 [15] 9 20221012A 45, 46 Bin2 0.15 11.81 9.77 0.2847 190 19.14 [15] 10 20171020A 69, 65 Bin3 0.09 9.69 0.04 0.0087 83 14.97 [47] 11 20190714A 51, 50 Bin2 1.89 9.94 1.66 0.2365 335 20.34 [5] 12 20220207C 80, 74 Bin4 1.16 9.85 0.81 0.0430 149 18.24 [15] 13 20220509G 60, 57 Bin3 0.23 11.43 6.17 0.0894 154 17.73 [15] 14 20220825A 46, 46 Bin2 1.62 9.74 0.89 0.2414 445 20.03 [15] 15 20220912A 43, 46 Bin2 2.5 9.6 1.0 0.077 34 19.65 [48] 16 20220307B 44, 47 Bin2 2.71 9.61 1.1 0.2481 126 19.93 [15] 17 20180924B 34, 36 Bin1 0.62 10.6 2.45 0.3212 80 20.33 [5] 18 20190102C 50, 50 Bin2 0.4 10.09 0.49 0.2909 94 20.77 [5] 19 20190608B 37, 36 Bin1 7.03 9.71 3.63 0.1178 238 17.41 [5] 20 20191001A 53, 56 Bin2 18.28 9.47 5.37 0.2342 340 18.36 [5] 21 20220319D 25, 37 Bin1 0.21 10.84 1.45 0.0112 16 15.39 [15] 22 20210807D 41, 41 Bin1 0.63 11.17 9.33 0.1293 66 17.17 [5] 23 20210320C 28, 27 Bin1 3.51 9.82 2.34 0.2796 148 19.47 [5]
a: Boldface indicates FRB hosts for which Pan-STARRS DR2 r-band images were used. For all other hosts, we use DESI DR10 r-band images.
b: The first inclination angle value in each row is computed using parametric profile fitting, and the second one is computed with the elliptical isophote fitting. The 1 error in both the estimates is .
c: SFR is the integrated 0–100 Myr star-formation rate taken from the references stated in the last column.
Name | Inclination angle rangea |
Degrees | |
Bin 1 | [0 42) |
Bin 2 | [42 60) |
Bin 3 | [60 76) |
Bin 4 | [76 90] |
a: Parentheses ‘)’ indicate that the endpoint is not included in the range, while square brackets ‘]’ indicate that the endpoint is included.